# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v2

@@ 4.4 Black Hole Shadow
This subsection discusses the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM. The critical shadow radius of a black hole is determined by the unstable photon sphere, which can be calculated from the effective potential of photons moving in gravitational field, or through a geometric approach proposed in a recent works [111]. Here, we calculate the radius of unstable photon sphere and the black hole shadow radius based on the geometric approach, in which the photon sphere and its stability are uniquely determined from Gaussian curvature and geodesic curvature in the optical geometry of black hole spacetime. Firstly, any photon orbits maintain their geodesic nature when they are transformed into optical geometry, so the geodesic curvature for a photon sphere vanishes. This geodesic curvature condition for photon sphere is completely equivalent to find the local extreme point of effective potential [111, 112, 113]
$\kappa_{g}(r=r_{\text{ph}})=0\ \ \Leftrightarrow\ \ \frac{dV_{\text{eff}}(r)}{dr}\bigg{|}_{r=r_{\text{ph}}}=0$ (28)
$\mathcal{K}(r=r_{\text{unstable}})<0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{unstable}}}<0$ (29a)
$\mathcal{K}(r=r_{\text{stable}})>0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{stable}}}>0$ (29b)
The absolute apparent size of black hole shadow viewed by an observer located at infinity (the asymptotically flat region) is characterized by critical shadow radius. With the unstable photon sphere obtained in equations (28) and (29a), the critical shadow radius of a black hole can be calculated through the relation
$r_{sh}=b_{\text{critical}}=\frac{r_{\text{unstable}}}{\sqrt{f(r=r_{\text{unstable}})}}$ (30)
@@
The numerical results on black hole shadow for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 6. The left panels plot the critical shadow radius $r_{sh}$ calculated using expression (30), and the right panels illustrate the black hole shadow in celestial coordinates, highlighting the black hole shadow affected by three black hole parameters (the black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$). From figure 6, it is reasonable to draw the following conclusions. Firstly, the increasing of black hole electric charge diminishes the apparent size of black hole shadow. Secondly, the lower left panel shows that the nonlinear electrodynamics parameter $a$ has very tiny influences on the black hole shadow size. For the same electric charge and PFDM parameter, the critical shadow radius calculated with different nonlinear electrodynamics parameter values do not exhibit any obvious differences. Thirdly, when PFDM parameter is not vary large, such as $\lambda_{\text{DM}}\leq 0.5M$ in upper and middle panels of figure 6, the size of black hole shadow is enlarged as PFDM parameter reduced. However, this statement is no longer valid when PFDM reaches a large critical point, see figure 7.
@@
To further explore the dark matter effects on black hole shadow size, we use figure 7 to show the critical shadow radius changed with respect to PFDM parameter, both in the presence and absence of nonlinear electrodynamics effects. It can be observed that when PFDM parameter is smaller than a critical value ($\lambda_{\text{DM}}<\lambda_{\text{critical}}$), the black hole shadow radius reduces with the increasing of PFDM parameter. However, when PFDM parameter excesses this critical value ($\lambda_{\text{DM}}>\lambda_{\text{critical}}$), the black hole shadow radius get magnified as the increasing of PFDM parameter. This seems to be a universal properties of the PFDM effects on black hole shadows. Recent studies on charged RN black hole surrounded by PFDM have reported similar conclusions as we have presented here (the existence of critical point $\lambda_{\text{critical}}$) [50, 49, 51]. Furthermore, figure 7 shows that the critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values. Since the nonlinear electrodynamics effects have tiny influences on black hole shadow size, the left and right panels of figure 7 exhibit very similar results and tends.
@@
Interestingly, the black hole shadow results for Euler-Heisenberg black hole surrounded by PFDM in the presence and absence of nonlinear electrodynamics effects / QED effects could also exhibit a slight difference. In the absence of nonlinear electrodynamics effects, because of the emergence of naked singularity when black hole charge $Q>Q_{\text{extreme}}$, the black hole shadow disappears for large black hole electric charge. On the other hand, in the presence of nonlinear electrodynamics effects, although the Euler-Heisenberg black hole (surrounded by PFDM) do not produce any naked singularities (see Appendix A for the discussions on horizons), the black hole shadow may still disappear under some circumstances. In the middle left panel of figure 6, when PFDM parameter takes the values $\lambda_{\text{DM}}=0.0M,0.1M,0.2M$, the critical shadow radius calculated using expression (30) is not continuous when black hole charge reaches a critical value $Q=Q_{\text{discontinuity}}$. This is caused by the discontinuity of unstable photon sphere positions in the vicinity of Euler-Heisenberg black hole. When black hole charge excesses this critical value, the number of unstable photon sphere around Euler-Heisenberg black hole undergoes a sudden change  [^3] When black hole charge satisfies $Q<Q_{\text{discontinuity}}$, there are two unstable photon spheres near the Euler-Heisenberg black hole (surrounded by PFDM), and the outer branch of unstable photon sphere determines the black hole shadow size via expression (30). When black hole charge becomes $Q>Q_{\text{discontinuity}}$, there is only one unstable photon sphere. The outer branch of unstable photon spheres in the $Q<Q_{\text{discontinuity}}$ cases is disappeared, which makes the unstable photon sphere position undergo a sudden change when black hole charge come across the critical point $Q=Q_{\text{discontinuity}}$.. Once the black hole charge comes across the discontinuity point (namely $Q>Q_{\text{discontinuity}}$), it is questionable that the results obtained using expression (30) could still represent a conventional “black hole shadow” or not. This issue deserves a more deeply study in the future.
@@ 5 Conclusions
In this work, the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM is studied. This kind of black hole solution, which are solved from the gravitational field equation coupled with nonlinear electromagnetic field and dark matter field, could give rise to the nonlinear electrodynamics effects and PFDM effects in charge black hole systems. It provide us a simply way to investigate the interplay of nonlinear electrodynamics effects (or QED effects) and dark matter effects on the studies of black holes. These effects coming from nonlinear electrodynamics and dark matter can be parameterized by parameters $a$ and $\lambda_{\text{DM}}$ respectively. In the presented work, we mainly focus on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angle of light, time delay of light, precession angle and trajectories of massive object’s bound orbits, black hole shadow are calculated numerically from geodesics.
@@
Our numerical results presented in this work show that influences coming from black hole electric charge, nonlinear electromagnetic effects, dark matter effects exhibit a similar tend for various gravitational lensing observables (including the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits, and the black hole shadow size). Firstly, the Euler-Heisenberg black hole surround by PFDM with a larger electric charge could reduce the gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius. Secondly, the nonlinear electrodynamics effects parameterized by $a$ has very tiny influences on these gravitational lensing observables. Thirdly, the most significant impacts on these gravitational lensing observables coming from the dark matter effects. The gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius are all greatly changed when PFDM parameter $\lambda_{\text{DM}}$ is varied. Particularly, a larger PFDM parameter could greatly reduce the gravitational deflection angle, time delay of light and precession angle of bound orbits. This can be explained from the spacetime metric function (6) for Euler-Heisenberg black hole surrounded by PFDM, in which the PFDM parameter plays a role of “effective mass”. A positive PFDM parameter contributes to a negative “effective mass” in the spacetime metric, therefore it can diminish the gravitational lensing observables. Moreover, the dark matter also has notable influences on the shape and eccentricity of massive object’s bound orbits. For black hole shadow size, when PFDM parameter is smaller than a critical value, the black hole shadow radius reduced as the increasing of PFDM parameter. When PFDM parameter exceeds this critical value, the black hole shadow radius could be magnified when PFDM parameter increases. The critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values.
@@
We hope that the results in our present work can deepen our understanding of the nonlinear electrodynamics effects / QED effects and dark matter effects on the gravitational lensing of charged black hole systems. Hopefully it may have potential applications for theoretical and observational probing for dark matter effects in black holes and gravitation. Since the dark matter could significantly influence the gravitational lensing observables, the PFDM parameter can be constrained from gravitational lensing observations, by analyzing and comparing the observed data with theoretical predictions, especially with the help of accumulating observational data for supermassive black hole in galaxy centers. However, the nonlinear electrodynamics effects have very tiny influences on these gravitational lensing observables, so it would be not easy to probe the nonlinear electrodynamics in gravitational lensing observations (unless very high-precision observational data are employed).
@@ Acknowledgment
The authors thanks Song-Lin Lyu and Guang-Zhou Guo for helpful discussion and comments on the numerical schemes. This research was funded by the Natural Science Foundation of Chongqing Municipality (Grant No. CSTB2022NSCQ-MSX0932), the Scientific and Technological Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202201126 and No. KJQN202301164), the Scientific Research Program of Chongqing Science and Technology Commission (the Chongqing “zhitongche” program for doctors, Grant No. CSTB2022BSXM-JCX0100), the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2020ZDZ027) and the Research and Innovation Team Cultivation Program of Chongqing University of Technology (Grant No. 2023TDZ007).

### v3

@@ 4.4 Black Hole Shadow
This subsection discusses the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM. The numerical results on black hole shadow for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 6. The left panels plot the critical shadow radius $r_{sh}$ calculated using expression (30), and the right panels illustrate the black hole shadow in celestial coordinates, highlighting the black hole shadow affected by three black hole parameters (the black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$). From figure 6, it is reasonable to draw the following conclusions. Firstly, the increasing of black hole electric charge diminishes the apparent size of black hole shadow. Secondly, the lower left panel shows that the nonlinear electrodynamics parameter $a$ has very tiny influences on the black hole shadow size. For the same electric charge and PFDM parameter, the critical shadow radius calculated with different nonlinear electrodynamics parameter values do not exhibit any obvious differences. Thirdly, when PFDM parameter is not vary large, such as $\lambda_{\text{DM}}\leq 0.5M$ in upper and middle panels of figure 6, the size of black hole shadow is enlarged as PFDM parameter reduces. However, this statement is no longer valid when PFDM reaches a large critical point, see figure 7.
@@
To further explore the dark matter effects on black hole shadow size, we use figure 7 to show the critical shadow radius changed with respect to PFDM parameter, both in the presence and absence of nonlinear electrodynamics effects. It can be observed that when PFDM parameter is smaller than a critical value ($\lambda_{\text{DM}}<\lambda_{\text{critical}}$), the black hole shadow radius reduces with the increasing of PFDM parameter. However, when PFDM parameter exceeds this critical value ($\lambda_{\text{DM}}>\lambda_{\text{critical}}$), the black hole shadow radius gets magnified as the increasing of PFDM parameter. This seems to be a universal property of the PFDM effects on black hole shadows. Recent studies on charged RN black hole surrounded by PFDM have reported similar conclusions as what we have presented here (the existence of critical point $\lambda_{\text{critical}}$) [51, 50, 52]. Furthermore, figure 7 shows that the critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values. Since the nonlinear electrodynamics effects have tiny influences on black hole shadow size, the left and right panels of figure 7 exhibit very similar results and tendency.
@@
Interestingly, the black hole shadow results for Euler-Heisenberg black hole surrounded by PFDM in the presence and absence of nonlinear electrodynamics effects / QED effects could also exhibit a slight difference. In the absence of nonlinear electrodynamics effects, because of the emergence of naked singularity when black hole charge $Q>Q_{\text{extreme}}$, the black hole shadow disappears for large black hole electric charge. On the other hand, in the presence of nonlinear electrodynamics effects, although the Euler-Heisenberg black hole (surrounded by PFDM) does not produce any naked singularities (see Appendix B for the discussions on horizons), the black hole shadow may still disappear under some circumstances. In the middle left panel of figure 6, when PFDM parameter takes the values $\lambda_{\text{DM}}=0.0M,0.1M,0.2M$, the critical shadow radius calculated using expression (30) is not continuous when black hole charge reaches a critical value $Q=Q_{\text{discontinuity}}$. This is caused by the discontinuity of unstable photon sphere positions in the vicinity of Euler-Heisenberg black hole. When black hole charge exceeds this critical value, the number of unstable photon sphere around Euler-Heisenberg black hole undergoes a sudden change  [^3] When black hole charge satisfies $Q<Q_{\text{discontinuity}}$, there are two unstable photon spheres near the Euler-Heisenberg black hole (surrounded by PFDM), and the outer branch of unstable photon sphere determines the black hole shadow size via expression (30). When black hole charge becomes $Q>Q_{\text{discontinuity}}$, there is only one unstable photon sphere. The outer branch of unstable photon spheres in the $Q<Q_{\text{discontinuity}}$ cases is disappeared, which makes the unstable photon sphere position undergo a sudden change when black hole charge passes the critical point $Q=Q_{\text{discontinuity}}$.. Once the black hole charge comes across the discontinuity point (namely $Q>Q_{\text{discontinuity}}$), it is questionable that the results obtained using expression (30) could still represent a conventional “black hole shadow” or not. This issue deserves a more deeply study in the future.
@@ 5 Conclusions
In this work, the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM is studied. This kind of black hole solution, which is solved from the gravitational field equation coupled with nonlinear electromagnetic field and dark matter field, could give rise to the nonlinear electrodynamics effects and PFDM effects in charge black hole systems. It provides us a simple way to investigate the interplay of nonlinear electrodynamics effects (or QED effects) and dark matter effects on the studies of black holes. These effects coming from nonlinear electrodynamics and dark matter can be parameterized by parameters $a$ and $\lambda_{\text{DM}}$ respectively. In the presented work, we mainly focus on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angle of light, time delay of light, precession angle and trajectories of massive object’s bound orbits, black hole shadow are calculated numerically from geodesics.
@@
Our numerical results presented in this work show that influences coming from black hole electric charge, nonlinear electromagnetic effects, dark matter effects exhibit a similar tendency for gravitational lensing observables (such as gravitational deflection angle of light and time delay of light), precession angle of massive object’s bound orbits, and the black hole shadow size. Firstly, the Euler-Heisenberg black hole surround by PFDM with a larger electric charge could reduce the gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius. Secondly, the nonlinear electrodynamics effects parameterized by $a$ has very tiny influences on these quantities. Thirdly, the most significant impacts on these observables are coming from the dark matter effects. The gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius are all greatly changed when PFDM parameter $\lambda_{\text{DM}}$ is varied. Particularly, a larger PFDM parameter could greatly reduce the gravitational deflection angle, time delay of light and precession angle of bound orbits. This can be explained from the spacetime metric function (6) for Euler-Heisenberg black hole surrounded by PFDM, in which the PFDM parameter plays a role of “effective mass”. A positive PFDM parameter contributes to a negative “effective mass” in the spacetime metric, therefore it can diminish the gravitational lensing observables (such as gravitational deflection angle and time delay of light). Moreover, the dark matter also has notable influences on the shape and eccentricity of massive object’s bound orbits. For black hole shadow size, when PFDM parameter is smaller than a critical value, the black hole shadow radius gets reduced as the increasing of PFDM parameter. When PFDM parameter exceeds this critical value, the black hole shadow radius could be magnified when PFDM parameter increases. The critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values.
@@
We hope that the results in our present work can deepen our understanding of the nonlinear electrodynamics effects / QED effects and dark matter effects on the gravitational lensing of charged black hole systems. Hopefully, it may have potential applications for theoretical and observational probing for dark matter effects in black holes and gravitation. Since the dark matter could significantly influence the gravitational lensing of supermassive black holes, the PFDM parameter can be constrained from gravitational lensing observations, by analyzing and comparing the observed data with theoretical predictions, especially with the help of accumulating observational data for supermassive black hole in galaxy centers. However, the nonlinear electrodynamics effects have very tiny influences on gravitational lensing observables, so it would be not easy to probe the nonlinear electrodynamics in gravitational lensing observations (unless very high-precision observational data are employed).
@@ Acknowledgment
The authors thank Song-Lin Lyu and Guang-Zhou Guo for helpful discussion and comments on the numerical schemes. This research was funded by the Natural Science Foundation of Chongqing Municipality (Grant No. CSTB2022NSCQ-MSX0932), the Scientific and Technological Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202201126 and No. KJQN202301164), the Scientific Research Program of Chongqing Municipal Science and Technology Bureau (the Chongqing “zhitongche” program for doctors, Grant No. CSTB2022BSXM-JCX0100), the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2020ZDZ027) and the Research and Innovation Team Cultivation Program of Chongqing University of Technology (Grant No. 2023TDZ007).
Abbreviations
The following abbreviations are used in this manuscript:
QED Quantum Electrodynamics
PFDM Perfect Fluid Dark Matter
RN Reissner-Nordström
WIMPs Weakly Interacting Massive Particles
EHT Event Horizon Telescope